he thermite reaction, used for welding iron, is the reaction of Fe3O4 with Al. 8 Al (s) + 3 Fe3O4 (s) \longrightarrow ⟶ 4 Al2O3 (s) + 9 Fe (s) \Delta Δ H° = -3350. kJ/mol rxn Because this large amount of heat cannot be rapidly dissipated to the surroundings, the reacting mass may reach temperatures near 3000. °C. How much heat (in kJ) is released by the reaction of 16 g of Al with 76.3 g of Fe3O4? Enter a positive number since released already tells us it is a negative number (to 1 decimal place).
Answer #1

Given reaction:

8 Al (s) + 3 Fe3O4 (s) ⟶ 4 Al2O3 (s) + 9 Fe (s) ΔH° = -3350. kJ/mol

Mass of Al = 16g

Molar mass of Al = 26.98g/mol

Moles of Al = 16/26.98 = 0.6 mol

Mass of Fe3O4 = 76.3g

Molar mass of Fe3O4 = 231.6g/mol

Moles of Fe3O4 = 76.3/231.6 = 0.33 mol

From the first equation, the molar ratio between Al and Fe3O4 is :

8 : 3 = 1 : 0.375

But, from the above calculations, the molar ratio between Al and Fe3O4 is :

0.6 : 0.33 = 1 : 0.55

Thus, Fe3O4 is in excess and Al is the limiting reagent.

So, the reaction stopped after 0.6 moles of Al was used.

Given, ΔH° = -3350. kJ/mol

Heat released = (0.6*3350)/8

Heat released = 251.25 kJ